A dentist invested a portion of $14,000 in a 8% annual simple interest account and the remainder in a 6.5% annual simple interest government bond. The two investments earn $1045 in interest annually. How much was invested in each account? in the 8%_____ and in the 6.5%______
HELP ME!!! =o)))) PLEASE

Answer by jim_thompson5910(35256)

x = amount (in dollars) invested in the 8% account
y = amount (in dollars) invested in the 6.5% account


$14,000 was invested in only two accounts, which means x and y must add to 14,000

x+y = 14000
y = 14000-x

I solved for y so I could plug it into the next equation below

-------------------------------------------------------

Now set up the second equation and plug in y = 14000-x. Then solve for x

0.08x + 0.065y = 1045
0.08x + 0.065(14000-x) = 1045
0.08x + 0.065(14000)-0.065(x) = 1045
0.08x + 910-0.065x = 1045
0.015x + 910 = 1045
0.015x = 1045 - 910
0.015x = 135
x = 135/0.015
x = 9000

Use this to find y

y = 14000-x
y = 14000-9000
y = 5000

How much was invested in each account?

in the 8% _____
and in the 6.5%______
------------------------------------------------------------------------------------------------------------------------

If you need more one-on-one help, email me at jim_thompson5910@hotmail.com. You can ask me a few more questions for free, but afterwards, I would charge you ($2 a problem to have steps shown or $1 a problem for answer only).

Alternatively, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Any amount is greatly appreciated as it helps me a lot. This donation is to support free tutoring. Thank you.

Jim
------------------------------------------------------------------------------------------------------------------------